OpenStudy (anonymous):

6x-30/24 how do i simplify this rational expression

OpenStudy (kropot72):

Take out the common factor 6.\[6x-\frac{30}{24}=6(x-\frac{5}{24})\]

OpenStudy (anonymous):

sorry its (6x-30)/24

OpenStudy (kropot72):

Divide each term in the numerator by 24\[\frac{6x-30}{24}=\frac{6x}{24}-\frac{30}{24}=\frac{x}{4}-1.25\]

OpenStudy (anonymous):

o seems simple enough

OpenStudy (kropot72):

Gr8!